How much pure acid should be mixed with 5 gallons of a 40% acid solution in order to get a 90% acid solution?
Answer by stanbon(75887)

How much pure acid should be mixed with 5 gallons of a 40% acid solution in order to get a 90% acid solution?
----
Equation:
acid + acid = acid
x + 0.40*5 = 0.90(x+5)
----------------------------
Multiply thru by 100 to get:
100x + 40*5 = 90x + 90*5
10x = 50*5
x = 25 gallons (amt. of pure acid needed for the mixture)
============================================
Cheers,
Stan H.